Can you please help!! My wife and I have been out of school for more than 15 years and we are home schooling and need help solving this problem. Could you let us know how to get the answer step by step please.
Paul is 5 years older than twice Carlo's age. The sum of their age is 101. How old is Carlo's.

Answer by usyim88hk(158)

First, let x be the age of Carlo
So x is the age of Carlo and 2x+5 is the age of Paul because it is twice the age of Carlo and plus five
Now we knwo that the sum of their age is 101, so we could set up an equation:
:
x + 2x + 5 = 101
(Age of Carlo plus age of Paul is equal to 101)
3x + 5 = 101
3x = 96
x = 32
So the age of Carlo is 32 and the age of Paul is 32 times 2 and plus 5 which is equal to 69